OpenStudy (anonymous):

A spinner is divided into 10 equal sections numbered 1 through 10. If the arrow is spun twice, what is the probability the first number will be a 2 and the second number will be a 4?

OpenStudy (lgbasallote):

i'm thinking \[\large \frac{1}{10} \times \frac{1}{10}\]

OpenStudy (anonymous):

0.01

OpenStudy (lgbasallote):

seems right